# Extended-Use Designs on Very Large Online Platforms

### v1

@@ 7.4. A Lack of Autonomy-Supporting Experiences For Teens
Perhaps one of the most notable aspects of our dataset of the EUDs a teen would encounter on VLOPs was the lack of teen-specific experiences. The designs platforms employed when a minor user was on the platform were consistent with those an adult might expect to encounter. Children and teens will be particularly susceptible to EUDs due to higher sensitivity to social feedback from peers, which will make pressure more effective in prolonging engagement. The behavior of children and teens is also more easily shaped through reinforcement schedules and pleasure-seeking, which makes them more susceptible to enticement. With less developed critical thinking skills, meta-cognition, and self-monitoring, children and teens are more likely to follow the flow of navigation cues and be trapped or lulled by deceptive designs. It is notable that these design categories frequently appear in the digital products that children are most likely to use through school hours and overnight—namely, social media, video-sharing, and gaming platforms (Monge Roffarello et al., 2023).
Here, we envision several possible design approaches that address the problematic aspects of EUDs (see Table 2). Because EUDs vary in their degree of manipulation, age-inappropriateness, and ethics, we propose solutions that range from prohibiting certain designs on platforms used by minors, to providing low-friction options to turn off a given design pattern. These counter-designs serve the purpose of increasing user agency and transparency to child and teen users about their options. Shifting design patterns from promoting extended use to supporting self-determination and disengagement would support youth wellbeing by reducing excessive time online and improving families’ sense of self-efficacy around technology.
@@
Table 2. Potential design approaches that combat EUDs, support users’ autonomy, and protect minors.
@@ 7.5. Limitations
There were several limitations to our ability to review VLOP interfaces. First, we did not create paid membership accounts, which prevented us from capturing features like Amazon Prime’s one-click-buy and other experiences designed for premium users. Second, we only examined Android and Windows systems, as these are the most-used operating systems in Europe, and VLOPs on iOS systems have slight differences in their design. Third, our short-term observation method cannot capture EUDs that arise over prolonged use, nor does it fully emulate the experience of a teenager interacting with the platform. We simulated the usage patterns of a teenage user by creating youth personas and accounts, which we found to yield valuable insights, but future work with actual teenage users would be a valuable complement to the data presented here.
Finally, this research was conducted outside of a corporate context. EUDs that are not readily observable (e.g., data tracking, personalization, and backend designs) would not have been possible for us to detect, and many of the proposed interventions would be difficult or impossible to implement without the cooperation of VLOPs themselves. Greater transparency and collaboration between researchers and industry stakeholders are crucial for building on the findings we present.
@@ 8. Conclusion
We conducted an inductive-deductive thematic analysis of all VLOPs recognized by the European Union to identify and characterize the EUDs these platforms use with teen users. We then distilled these into a taxonomy and found that designs clustered into four categories reflecting the strategies that underlie them: 1) pressuring, 2) enticing, 3) trapping, and 4) lulling users into engaging. These approaches manipulate users’ emotional experience (i.e., pressuring, enticing) and users’ cognitive load (i.e., lulling, trapping). They work together to reward users emotionally and cognitively for continuing to engage and to punish users emotionally and cognitively for attempting to disengage.
We then conducted a structured content analysis, applying our taxonomy to all 17 VLOPs. This revealed that EUDs are pervasive, with 583 features employing these approaches. Although all VLOPs used at least some EUDs, social media platforms used twice as many as other VLOPs. In the wild, EUDs do not operate in isolation, but are chained together to trigger one another, act simultaneously on the user, and combine into single features, thus turning what might, in isolation, be subtle nudges into a cacophony of aggressive attention-grabbing gimmicks.

### v2

@@ 7.4. A Lack of Autonomy-Supporting Experiences For Teens
Perhaps one of the most notable aspects of our dataset of the EPDs a teen would encounter on VLOPs was the lack of teen-specific experiences. The designs platforms employed when a minor user was on the platform were consistent with those an adult might expect to encounter. Children and teens will be particularly susceptible to EPDs due to higher sensitivity to social feedback from peers, which will make pressure more effective in prolonging engagement. The behavior of children and teens is also more easily shaped through reinforcement schedules and pleasure-seeking, which makes them more susceptible to enticement. With less developed critical thinking skills, meta-cognition, and self-monitoring, children and teens are more likely to follow the flow of navigation cues and be trapped or lulled by deceptive designs. It is notable that these design categories frequently appear in the digital products that children are most likely to use through school hours and overnight—namely, social media, video-sharing, and gaming platforms (Monge Roffarello et al., 2023).
Here, we envision several possible design approaches that address the problematic aspects of EPDs (see Table 2). Because EPDs vary in their degree of manipulation, age-inappropriateness, and ethics, we propose solutions that range from prohibiting certain designs on platforms used by minors, to providing low-friction options to turn off a given design pattern. These counter-designs serve the purpose of increasing user agency and transparency to child and teen users about their options. Shifting design patterns from promoting extended use to supporting self-determination and disengagement would support youth wellbeing by reducing excessive time online and improving families’ sense of self-efficacy around technology.
@@
Table 2. Potential design approaches that combat EPDs, support users’ autonomy, and protect minors.
@@ 7.5. Limitations
There were several limitations to our ability to review VLOP interfaces. First, we did not create paid membership accounts, which prevented us from capturing features like Amazon Prime’s one-click-buy and other experiences designed for premium users. Second, we only examined Android and Windows systems, as these are the most-used operating systems in Europe, and VLOPs on iOS systems have slight differences in their design. Third, our short-term observation method cannot capture EPDs that arise over prolonged use, nor does it fully emulate the experience of a teenager interacting with the platform. We simulated the usage patterns of a teenage user by creating youth personas and accounts, which we found to yield valuable insights, but future work with actual teenage users would be a valuable complement to the data presented here.
Finally, this research was conducted outside of a corporate context. EPDs that are not readily observable (e.g., data tracking, personalization, and backend designs) would not have been possible for us to detect, and many of the proposed interventions would be difficult or impossible to implement without the cooperation of VLOPs themselves. Greater transparency and collaboration between researchers and industry stakeholders are crucial for building on the findings we present.
@@ 8. Conclusion
We conducted an inductive-deductive thematic analysis of all VLOPs recognized by the European Union to identify and characterize the EPDs these platforms use with teen users. We then distilled these into a taxonomy and found that designs clustered into four categories reflecting the strategies that underlie them: 1) pressuring, 2) enticing, 3) trapping, and 4) lulling users into engaging. These approaches manipulate users’ emotional experience (i.e., pressuring, enticing) and users’ cognitive load (i.e., lulling, trapping). They work together to reward users emotionally and cognitively for continuing to engage and to punish users emotionally and cognitively for attempting to disengage.
We then conducted a structured content analysis, applying our taxonomy to all 17 VLOPs. This revealed that EPDs are pervasive, with 583 features employing these approaches. Although all VLOPs used at least some EPDs, social media platforms used twice as many as other VLOPs. In the wild, EPDs do not operate in isolation, but are chained together to trigger one another, act simultaneously on the user, and combine into single features, thus turning what might, in isolation, be subtle nudges into a cacophony of aggressive attention-grabbing gimmicks.
Selection and Participation of Children
No children participated in this work.